OpenStudy (anonymous):

Which value of a solves the equation? a-2=1/64

OpenStudy (anonymous):

\[a-\frac{128}{64}=\frac{1}{64}, a= \frac{129}{64}, a= 2\frac{1}{64}\]

OpenStudy (anonymous):

Um, the 2 is small. The options are A. –32 B. –16 C. 8 D. 4

OpenStudy (anonymous):

Did you type it right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[a-2=\frac{1}{64}?\]